# appendix,by

### v1

@@ 4.2 Weak quantifier elimination and weak stable embeddedness
In this subsection we generalise to positive logic certain lemmas from [CH14] concerning array extension.
Recall that usual inductive theories may be identified with Pillay positive theories; see [DM24, Subsection 6.2] for details.
@@
This is a standard application of a general topological fact, see e.g. [TZ12, Exercise 3.1.1]. In the notation of the latter, given $\varphi(x)$ a quantifier-free formula, take as $X$ the space of maximal types in variables $x$, as $Y_{1}$ the set given by $\varphi(x)$, which is clopen because $T$ is Pillay, and as $Y_{2}$ its complement. ∎
@@
Let $)\lambda=\beth((2^{\lvert T\rvert+\lvert abc_{00}\rvert}$$^{+})$. By compactness, we can extend $(c_{ij})_{i,j<\omega}$ to a strongly indiscernible array $(c_{ij})_{i<\lambda,j<\omega}$ such that $\bar{c}_{<\lambda}$ is $a$-indiscernible. For all $i<\lambda$ and $j<\omega$, let $b^{\prime}_{ij}$ be such that $b^{\prime}_{ij}c_{ij}\equiv b_{00}c_{00}$. By ˜4.10 we can find a strongly indiscernible array $(b^{\prime\prime}_{ij}c^{\prime\prime}_{ij})_{i<\lambda,j<\omega}$ which is $\mathrm{EM}_{\mathrm{ar}}$-based on $(\bar{b}^{\prime}_{i}\bar{c}_{i})_{i<\lambda}$. As $\bar{c}_{<\lambda}$ was strongly indiscernible, $\mathrm{EM}_{\mathrm{ar}}((\bar{b}^{\prime}_{i}\bar{c}_{i})_{i<\lambda})\supseteq\operatorname{tp}(\bar{c}_{<\lambda}$ so $\bar{c}^{\prime\prime}_{<\lambda}\equiv\bar{c}_{<\lambda}$. Hence, by applying an automorphism, we may assume $\bar{c}^{\prime\prime}_{<\lambda}=\bar{c}_{<\lambda}$. By ˜4.3, we can find an $a$-indiscernible sequence $(\bar{b}_{i}\bar{c}^{\prime\prime\prime}_{i})_{i<\omega}$ based on $(\bar{b}^{\prime\prime}_{i}\bar{c}_{i})_{i<\lambda}$ over $a$. As $\bar{c}_{<\lambda}$ was already indiscernible over $a$, we get that $\bar{c}^{\prime\prime\prime}_{<\omega}\equiv_{a}\bar{c}_{<\omega}$, so, by applying an automorphism over $a$, we may assume $\bar{c}^{\prime\prime\prime}_{<\omega}=\bar{c}_{<\omega}$. Then $(b_{ij})_{i,j<\omega}$ satisfies the requirements. ∎
@@
We may assume that $C=\emptyset$. By Proposition˜4.9 there is a mutually indiscernible array $(\bar{a}^{\prime\prime}_{i})_{i<\lambda}$ which is $\mathrm{EM}_{\mathcal{L}_{\lambda}}$–based on $\bar{a}_{<\lambda}$. By the assumption on $\bar{a}_{<\lambda}$ we have that $\bar{a}_{i}$ is indiscernible for every $i<\lambda$, hence $\bar{a}^{\prime\prime}_{i}\equiv\bar{a}_{i}$, and that for all $k<\omega$, $i_{0}<\ldots<i_{k-1}<\lambda$ and $j_{0},\ldots,j_{k-1}<\omega$ we have $a_{i_{0}j_{0}}\ldots a_{i_{k-1}j_{k-1}}\equiv_{C}a_{i_{0}0}\ldots a_{i_{k-1}0}$, hence $(a^{\prime\prime}_{i0})_{i<\lambda}\equiv(a_{i0})_{i<\lambda}$. Thus, we can find an automorphism $f$ sending $(a^{\prime\prime}_{i0})_{i<\lambda}$ to $(a_{i0})_{i<\lambda}$ and define $\bar{a}^{\prime}_{i}=f(\bar{a}^{\prime\prime}_{i})$ for each $i<\lambda$. ∎
@@
Then $b^{+}_{i}c^{+}_{i0}\equiv_{a}bc_{00}$ for every $i<\lambda$, and, by the inductive assumption and ˜4.30(1) applied to $k$ times, one for each row of $(\bar{b}^{*}_{i}\bar{c}^{*}_{i})_{i<k}$,
@@ 4.3 $\mathsf{NTP}_{2}$ for existentially closed valued difference fields of residual characteristic zero
In our proof of $\mathsf{NTP}_{2}$ for valued difference fields of residual characteristic zero we will need to Morleyise the residue field sort. For details on (partial) Morleyisation see [DK22, Remark 2.2] and [DM24, proof of Remark 6.14].
@@
In this subsection we denote by $\mathcal{L}_{0}$ one of the four languages in ˜3.22. We let $\mathrm{Mor}_{k}$ be the language of the Morleyisation of $\mathsf{ACFA}$. We write $\mathcal{L}$ for the expansion of $\mathcal{L}_{0}$ by the symbols in $\mathrm{Mor}_{k}$, added on the $k$ sort. We write $\mathsf{ACVFA}_{0,0,\mathcal{L}}$ for the h-inductive theory of $\mathcal{L}$-valued difference fields of residual characteristic zero. By $\mathsf{ACVFA}_{0,0}$ we mean the h-inductive theory of valued difference fields of residual characteristic zero in the language $\mathcal{L}_{\sigma}$ of ˜1.4.
Note that, contrary to $\mathsf{ACFA}$, the axioms of $\mathsf{ACVFA}_{0,0,\mathcal{L}}$ and of $\mathsf{ACVFA}_{0,0}$ do not require any form of genericity of $\sigma$. This is common in the literature (e.g. [HK21, DM24]), and due to the fact that, at any rate, we focus on the existentially closed models.
@@
$\mathsf{ACVFA}_{0,0,\mathcal{L}}$ is a Pillay theory, as the negations of the formulas $x<_{\Gamma}y$, $x=_{\Gamma}y$, and $x=_{\mathrm{VF}}y$ are equivalent to the positive formulas $x>_{\Gamma}y\vee x=_{\Gamma}y$, $x<_{\Gamma}y\vee x>_{\Gamma}y$, and $(x-y)^{-1}(x-y)=_{\mathrm{VF}}1$, respectively, and the negation of every predicate coming from Morleyising the sort $k$ is also added to the language in the Morleyisation process.
@@
The JCP-refinements of $\mathsf{ACVFA}_{0,0,\mathcal{L}}$ and of $\mathsf{ACVFA}_{0,0}$ are determined by fixing a characteristic 0 completion of $\mathsf{ACFA}$.
@@
Indeed, suppose $T$ is such a completion and $A,B\vDash\mathsf{ACVFA}_{0,0,\mathcal{L}}\cup T$. Then the prime subfields of $k(A)$ and of $k(B)$ are $\mathrm{Mor}_{k}$-isomorphic. By identifying both of them with $\mathbb{Q}$, we get that $k(A)$ and $k(B)$ amalgamate over $\mathbb{Q}$, so, by ˜3.28, $A$ and $B$ amalgamate over $\mathbb{Q}$. This proves JCP.
Conversely, if $T\supseteq\mathsf{ACVFA}_{0,0,\mathcal{L}}$ is a JCP-refinement, then for every $A,B\vDash T$ we can embed $A$ and $B$ in a model $C\vDash T$ and so the copies of $\mathbb{Q}^{\mathrm{alg}}$ in $k(A)$ and $k(B)$ are isomorphic. By [CH99, Proposition 1.4], this means that $T$ implies a completion of $\mathsf{ACFA}$ on $k$.
@@
$\mathsf{ACVFA}_{0,0,\mathcal{L}}$ admits weak quantifier elimination in $\mathcal{L}$.
@@
By [Hod93, Theorems 8.5.2 and 8.5.4], it suffices to show that the class of substructures of models of $\mathsf{ACVFA}_{0,0,\mathcal{L}}$ has the Amalgamation Property.
Let $L\leftarrow M\rightarrow N$ be substructures of $\mathcal{L}$-valued difference fields. By ˜3.28 it suffices to show that the amalgamation problem $k(L)\leftarrow k(M)\rightarrow k(N)$ has a solution. If $k_{1}$ and $k_{2}$ are characteristic 0 models of $\mathsf{ACFA}$ in the language $\mathrm{Mor}_{k}$ extending $k(L)$ and $k(N)$, respectively, such that $k_{2}$ is $\lvert k_{1}\rvert^{+}-$saturated, then $k(M)$ is a $\mathrm{Mor}_{k}$-substructure of $k_{1}$ and of $k_{2}$, hence $\mathrm{id}_{k(M)}$ is a partial $\mathrm{Mor}_{k}$-elementary map between $k_{1}$ and $k_{2}$. Hence $\mathrm{id}_{k(M)}$ extends to an embedding of $k_{1}$ into $k_{2}$. ∎
@@
In every existentially closed model of $\mathsf{ACVFA}_{0,0,\mathcal{L}}$ the sorts $\Gamma$ and $k$ are orthogonal, i.e. for every pair of maximal types $p(x)$ and $q(y)$ with $x$ a tuple of variables of the sort $\Gamma$ and $y$ a tuple of variables of the sort $k$, the partial type $p(x)\cup q(y)$ implies a maximal type in variables $xy$.
@@
In every existentially closed $\mathfrak{U}\vDash\mathsf{ACVFA}_{0,0,\mathcal{L}}$ the following hold.
@@
For (1), it is enough to show that if $M$ is a substructure of $\mathfrak{U}$ and $\gamma,\delta\in\Gamma(\mathfrak{U})^{n}$ are such that $\operatorname{qftp}(\gamma/\Gamma(M))=\operatorname{qftp}(\delta/\Gamma(M))$, then $\operatorname{tp}(\gamma/M)=\operatorname{tp}(\delta/M)$. Indeed, if we have this, let $\varphi(x,c)$ be a quantifier-free formula, with $x$ a tuple of variables of sort $\Gamma$, and set $M\coloneqq\operatorname{dcl}(c)$. Apply Lemma˜4.21 with $\mathcal{F}$ the family of quantifier-free formulas over $\Gamma(M)$, obtaining that $\varphi(x,c)$ is equivalent to such a formula. By Proposition˜4.35, it suffices to check the definition of strictly weakly stably embedded on quantifier-free formulas, hence this suffices.
Observe that, in the proofs of [Kes24, Lemmas 4.9 and 4.10], the assumption that $\sigma_{\Gamma}$ is multiplicative is never used. We apply [Kes24, Lemma 4.10] for the $\Gamma$ case, the other lemma being used in the $k$ case. By the cited result, we have an isomorphism $f\colon\langle M,\gamma\rangle\to\langle M,\delta\rangle$ extending $\mathrm{id}_{M}$ and sending $\gamma$ to $\delta$. As $k(\langle M,\gamma\rangle)=k(M)=k(\langle M,\delta\rangle)$, in particular, $f$ restricted to $k(\langle M,\gamma\rangle)$ is $\mathrm{Mor}_{k}$-elementary. Hence, by weak quantifier elimination we get that $f$ is a partial immersion, so in particular $\operatorname{tp}(\gamma/M)=\operatorname{tp}(\delta/M)$.
For (2) we show that, if $\gamma,\delta\in\Gamma(\mathfrak{U})^{n}$ and $\operatorname{qftp}_{\Gamma}(\gamma)=\operatorname{qftp}_{\Gamma}(\delta)$, then $\operatorname{tp}(\gamma)=\operatorname{tp}(\delta)$, and then apply Lemma˜4.21 and Proposition˜4.35. Again by [Kes24, Lemma 4.10], we have an isomorphism $\langle\mathbb{Q},\gamma\rangle\to\langle\mathbb{Q},\delta\rangle$ sending $\gamma$ to $\delta$. As $k(\langle\mathbb{Q},\gamma\rangle)=\mathbb{Q}=k(\langle\mathbb{Q},\delta\rangle)$, we have in particular that $f|_{k(\langle\mathbb{Q},\gamma\rangle)}=\mathrm{id}_{\mathbb{Q}}$ is elementary. Hence, by weak quantifier elimination we get that $f$ is a partial immersion.
For the other three languages, the proof is analogous, and uses analogues of [Kes24, Lemmas 4.9 and 4.10] that are easier to prove and left to the reader. ∎
@@
If $K\subseteq L$ is an immediate extension of $\mathcal{L}$-valued difference subfields of $\mathfrak{U}$, and $a\in L$, then every realisation in $\mathfrak{U}$ of the quantifier-free type of $a$ over $K$ in the language $\mathcal{L}_{0}$ (see ˜4.32) realises $\operatorname{tp}(a/K)$.
@@
This follows by Proposition˜4.35 and inspection of $\mathcal{L}$. ∎
@@
By ˜3.28, as observed in [DM24, Remark 6.55], the positive theory $\mathsf{ACVFA}_{0,0,\mathcal{L}}$ is Hausdorff, see [DK22, Definition 2.13]. In particular, it is thick.
@@
The positive theories $\mathsf{ACVFA}_{0,0,\mathcal{L}}$ and $\mathsf{ACVFA}_{0,0}$ are $\mathsf{NTP}_{2}$.
@@
If $M$ is an $\mathcal{L}$-valued difference field, or a valued difference field in the language $\mathcal{L}_{\sigma}$, as in the proof of Corollary˜3.10 we may expand a suitable pec extension of it to an $\mathcal{L}_{s,\iota,\sigma}$-valued difference field. It follows that it suffices to show that $\mathsf{ACVFA}_{0,0,s,\iota}$ is $\mathsf{NTP}_{2}$.
Suppose for a contradiction that a formula $\varphi(x,y)$ has $\mathsf{TP}_{2}$. We may assume that $x$ is a single variable by Corollary˜4.19. As the theory of $\Gamma$ is $\mathsf{NIP}$ (hence $\mathsf{NTP}_{2}$, see [DGK23, Theorem 1.1]) by [DM24, Corollary 6.39], by Proposition˜4.37(2) and ˜4.13 the structure $\Gamma_{\mathrm{ind}}(\mathfrak{U})$ is $\mathsf{NTP}_{2}$. Similarly for $k_{\mathrm{ind}}(\mathfrak{U})$, using the fact that $k(\mathfrak{U})\vDash\mathsf{ACFA}$, hence is simple [CH99]. Therefore, by Lemmas˜4.28 and 4.16 we may further assume that $x$ is of sort $\mathrm{VF}$. By Proposition˜4.35 and ˜4.13 we may furthermore assume that $\varphi(x,y)$ is quantifier-free.
Let $(a_{ij})_{i,j<\omega}$ be a strongly indiscernible array witnessing that $\varphi(x,y)$ has $\mathsf{TP}_{2}$. Let $a\vDash\{\varphi(x,a_{i0})\mid i<\omega\}$. By ˜4.4 we may assume that the sequence $(\bar{a}_{i})_{i<\omega}$ is $a$-indiscernible. We run the analogues for the language $\mathcal{L}_{s,\iota,\sigma,\mathrm{Mor}_{k}}$ of steps (I) to (V) in the proof of [CH14, Theorem 4.1], but we replace [CH14, Lemmas 3.6 and 3.8] by Lemmas˜4.31 and 4.26. We obtain, for $\beta<\omega$, strongly indiscernible arrays $(a_{ij}^{\beta})_{i,j<\omega}$ such that $K\coloneqq\bigcup_{\beta<\omega}a^{\beta}_{00}$ is a substructure of $\mathfrak{U}$ (in fact, a valued difference subfield because, since $\iota$ and $s$ are in the language, the maps $\operatorname{res}$ and $v$ are automatically surjective), the union is increasing, $L\coloneqq K\langle a\rangle$ is an immediate extension of $K$,

### v2

@@ 4.2 Weak quantifier elimination and weak stable embeddedness
In this subsection we introduce the notions of weak quantifier elimination and (strict) weak stable embeddedness, and generalise to positive logic certain lemmas from [CH] concerning array extension.
Recall that usual inductive theories may be identified with Pillay positive theories; see [DM, Subsection 6.2] for details.
@@
This is a standard application of a general topological fact, see e.g. [tentCourseModelTheory2012, Exercise 3.1.1]. In the notation of the cited source, given $\varphi(x)$ a quantifier-free formula, take as $X$ the space of maximal types in variables $x$, as $Y_{1}$ the set given by $\varphi(x)$, which is clopen because $T$ is Pillay, and as $Y_{2}$ its complement. ∎
@@
Let $)\lambda=\beth((2^{\lvert T\rvert+\lvert abc_{00}\rvert}$$^{+})$. By compactness, we can extend $(c_{ij})_{i,j<\omega}$ to a strongly indiscernible array $(c_{ij})_{i<\lambda,j<\omega}$ such that $\bar{c}_{<\lambda}$ is $a$-indiscernible. For all $i<\lambda$ and $j<\omega$, let $b^{\prime}_{ij}$ be such that $b^{\prime}_{ij}c_{ij}\equiv b_{00}c_{00}$. By ˜4.10 we can find a strongly indiscernible array $(b^{\prime\prime}_{ij}c^{\prime\prime}_{ij})_{i<\lambda,j<\omega}$ which is $\mathrm{EM}_{\mathrm{ar}}$-based on $(\bar{b}^{\prime}_{i}\bar{c}_{i})_{i<\lambda}$. As $\bar{c}_{<\lambda}$ was strongly indiscernible, $\mathrm{EM}_{\mathrm{ar}}((\bar{b}^{\prime}_{i}\bar{c}_{i})_{i<\lambda})\supseteq\operatorname{tp}(\bar{c}_{<\lambda})$ so $\bar{c}^{\prime\prime}_{<\lambda}\equiv\bar{c}_{<\lambda}$. Hence, by applying an automorphism, we may assume $\bar{c}^{\prime\prime}_{<\lambda}=\bar{c}_{<\lambda}$. By ˜4.3, we can find an $a$-indiscernible sequence $(\bar{b}_{i}\bar{c}^{\prime\prime\prime}_{i})_{i<\omega}$ based on $(\bar{b}^{\prime\prime}_{i}\bar{c}_{i})_{i<\lambda}$ over $a$. As $\bar{c}_{<\lambda}$ was already indiscernible over $a$, we get that $\bar{c}^{\prime\prime\prime}_{<\omega}\equiv_{a}\bar{c}_{<\omega}$, so, by applying an automorphism over $a$, we may assume $\bar{c}^{\prime\prime\prime}_{<\omega}=\bar{c}_{<\omega}$. Then $(b_{ij})_{i,j<\omega}$ satisfies the requirements. ∎
@@
We may assume that $C=\emptyset$. By Proposition˜4.9 there is a mutually indiscernible array $(\bar{a}^{\prime\prime}_{i})_{i<\lambda}$ which is $\mathrm{EM}_{\mathfrak{L}_{\lambda}}$–based on $\bar{a}_{<\lambda}$. By the assumption on $\bar{a}_{<\lambda}$ we have that $\bar{a}_{i}$ is indiscernible for every $i<\lambda$, hence $\bar{a}^{\prime\prime}_{i}\equiv\bar{a}_{i}$, and that for all $k<\omega$, $i_{0}<\ldots<i_{k-1}<\lambda$ and $j_{0},\ldots,j_{k-1}<\omega$ we have $a_{i_{0}j_{0}}\ldots a_{i_{k-1}j_{k-1}}\equiv_{C}a_{i_{0}0}\ldots a_{i_{k-1}0}$, hence $(a^{\prime\prime}_{i0})_{i<\lambda}\equiv(a_{i0})_{i<\lambda}$. Thus, we can find an automorphism $f$ sending $(a^{\prime\prime}_{i0})_{i<\lambda}$ to $(a_{i0})_{i<\lambda}$ and define $\bar{a}^{\prime}_{i}=f(\bar{a}^{\prime\prime}_{i})$ for each $i<\lambda$. ∎
@@
Then $b^{+}_{i}c^{+}_{i0}\equiv_{a}bc_{00}$ for every $i<\lambda$, and, by the inductive assumption and ˜4.30(1) applied $k$ times, one for each row of $(\bar{b}^{*}_{i}\bar{c}^{*}_{i})_{i<k}$,
@@ 4.3 $\mathsf{NTP}_{2}$ for existentially closed valued difference fields of residual characteristic zero
In our proof of $\mathsf{NTP}_{2}$ for valued difference fields of residual characteristic zero we will need to Morleyise the residue field sort. For details on (partial) Morleyisation in positive logic see [DK, Remark 2.2] and [DM, proof of Remark 6.14].
@@
In this subsection we denote by $\mathcal{L}$ one of the four languages in ˜3.22. We let $\mathrm{Mor}_{k}$ be the language of the Morleyisation of $\mathsf{ACFA}$. We write $\widehat{\mathcal{L}}$ for the expansion of $\mathcal{L}$ by the symbols in $\mathrm{Mor}_{k}$, added on the $k$ sort. We write $\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ for the h-inductive theory of $^\widehat{\mathcal{L}}$-valued difference fields of residual characteristic zero. By $\mathsf{ACVFA}_{0,0}$ we mean the h-inductive theory of valued difference fields of residual characteristic zero in the language $\mathcal{L}_{\sigma}$ of ˜1.4.
Note that, contrary to $\mathsf{ACFA}$, the axioms of $\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ and of $\mathsf{ACVFA}_{0,0}$ do not require any form of genericity of $\sigma$. This is common in the literature (e.g. [HK, DM]), and due to the fact that, at any rate, we focus on the existentially closed models.
@@
$\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ is a Pillay theory, as the negations of the formulas $x<_{\Gamma}y$, $x=_{\Gamma}y$, and $x=_{\mathrm{VF}}y$ are equivalent to the positive formulas $x>_{\Gamma}y\vee x=_{\Gamma}y$, $x<_{\Gamma}y\vee x>_{\Gamma}y$, and $(x-y)^{-1}(x-y)=_{\mathrm{VF}}1$, respectively, and the negation of every predicate coming from Morleyising the sort $k$ is also added to the language in the Morleyisation process.
@@
The JCP-refinements of $\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ and of $\mathsf{ACVFA}_{0,0}$ are determined by fixing a characteristic zero completion of $\mathsf{ACFA}$.
@@
Indeed, suppose $T$ is such a completion and $A,B\vDash\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}\cup T$. Then the prime subfields of $k(A)$ and of $k(B)$ are $\mathrm{Mor}_{k}$-isomorphic. By identifying both of them with the same $\mathrm{Mor}_{k}$-expansion of $\mathbb{Q}$, we get that $k(A)$ and $k(B)$ amalgamate over $\mathbb{Q}$, so, by ˜3.28, $A$ and $B$ amalgamate over $\mathbb{Q}$. This proves JCP.
Conversely, if $T\supseteq\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ is a JCP-refinement, then for every $A,B\vDash T$ we can embed $A$ and $B$ in a model $C\vDash T$ and so the copies of $\mathbb{Q}^{\mathrm{alg}}$ in $k(A)$ and $k(B)$ are isomorphic. By [CHr, Proposition 1.4], this means that $T$ implies a completion of $\mathsf{ACFA}$ on $k$.
@@
$\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ admits weak quantifier elimination in $\widehat{\mathcal{L}}$.
@@
By [Ho, Theorems 8.5.2 and 8.5.4], it suffices to show that the class of substructures of models of $\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ has the Amalgamation Property.
Let $L\leftarrow M\rightarrow N$ be substructures of $^\widehat{\mathcal{L}}$-valued difference fields. By ˜3.28 it suffices to show that the amalgamation problem $k(L)\leftarrow k(M)\rightarrow k(N)$ has a solution. If $k_{1}$ and $k_{2}$ are characteristic zero models of $\mathsf{ACFA}$ in the language $\mathrm{Mor}_{k}$ extending $k(L)$ and $k(N)$, respectively, such that $k_{2}$ is $\lvert k_{1}\rvert^{+}-$saturated, then $k(M)$ is a $\mathrm{Mor}_{k}$-substructure of $k_{1}$ and of $k_{2}$, hence $\mathrm{id}_{k(M)}$ is a partial $\mathrm{Mor}_{k}$-elementary map between $k_{1}$ and $k_{2}$. Hence $\mathrm{id}_{k(M)}$ extends to an embedding of $k_{1}$ into $k_{2}$. ∎
@@
In every existentially closed model of $\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ the sorts $\Gamma$ and $k$ are orthogonal, i.e. for every pair of maximal types $p(x)$ and $q(y)$ with $x$ a tuple of variables of the sort $\Gamma$ and $y$ a tuple of variables of the sort $k$, the partial type $p(x)\cup q(y)$ implies a maximal type in variables $xy$.
@@
In every existentially closed $\mathfrak{U}\vDash\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ the following hold.
@@
For (1), it is enough to show that if $M$ is a substructure of $\mathfrak{U}$ and $\gamma,\delta\in\Gamma(\mathfrak{U})^{n}$ are such that $\operatorname{qftp}(\gamma/\Gamma(M))=\operatorname{qftp}(\delta/\Gamma(M))$, then $\operatorname{tp}(\gamma/M)=\operatorname{tp}(\delta/M)$. Indeed, if we have this, let $\varphi(x,c)$ be a quantifier-free formula, with $x$ a tuple of variables of sort $\Gamma$, and set $M\coloneqq\operatorname{dcl}(c)$. Apply Lemma˜4.21 with $\mathcal{F}$ the family of quantifier-free formulas over $\Gamma(M)$, obtaining that $\varphi(x,c)$ is equivalent to such a formula. By Proposition˜4.35, it suffices to check the definition of strict weak stable embeddedness on quantifier-free formulas, hence this suffices.
Observe that, in the proofs of [Kesting, Lemmas 4.9 and 4.10], the assumption that $\sigma_{\Gamma}$ is multiplicative is never used. We apply [Kesting, Lemma 4.10] for the $\Gamma$ case, the other lemma being used in the $k$ case. By the cited result, we have an isomorphism $f\colon\langle M,\gamma\rangle\to\langle M,\delta\rangle$ extending $\mathrm{id}_{M}$ and sending $\gamma$ to $\delta$. As $k(\langle M,\gamma\rangle)=k(M)=k(\langle M,\delta\rangle)$, in particular, $f$ restricted to $k(\langle M,\gamma\rangle)$ is $\mathrm{Mor}_{k}$-elementary. Hence, by weak quantifier elimination we get that $f$ is a partial immersion, so in particular $\operatorname{tp}(\gamma/M)=\operatorname{tp}(\delta/M)$.
For (2) we show that, if $\gamma,\delta\in\Gamma(\mathfrak{U})^{n}$ and $\operatorname{qftp}_{\Gamma}(\gamma)=\operatorname{qftp}_{\Gamma}(\delta)$, then $\operatorname{tp}(\gamma)=\operatorname{tp}(\delta)$, and then apply Lemma˜4.21 and Proposition˜4.35. Again by [Kesting, Lemma 4.10], we have an isomorphism $\langle\mathbb{Q},\gamma\rangle\to\langle\mathbb{Q},\delta\rangle$ sending $\gamma$ to $\delta$. As $k(\langle\mathbb{Q},\gamma\rangle)=\mathbb{Q}=k(\langle\mathbb{Q},\delta\rangle)$, we have in particular that $f|_{k(\langle\mathbb{Q},\gamma\rangle)}=\mathrm{id}_{\mathbb{Q}}$ is elementary. Hence, by weak quantifier elimination we get that $f$ is a partial immersion.
For the other three languages, the proof is analogous, and uses analogues of [Kesting, Lemmas 4.9 and 4.10] that are easier to prove and left to the reader. ∎
@@
If $K\subseteq L$ is an immediate extension of $^\widehat{\mathcal{L}}$-valued difference subfields of $\mathfrak{U}$, and $a\in L$, then every realisation in $\mathfrak{U}$ of the quantifier-free type of $a$ over $K$ in the language $\mathcal{L}$ (see ˜4.32) realises $\operatorname{tp}(a/K)$.
@@
This follows by Proposition˜4.35 and inspection of $\widehat{\mathcal{L}}$. ∎
@@
By ˜3.28, as observed in [DM, Remark 6.55], the positive theory $\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ is Hausdorff, see [DK, Definition 2.13]. In particular, it is thick.
@@
The positive theories $\mathsf{ACVFA}_{0,0,\widehat{\mathcal{L}}}$ and $\mathsf{ACVFA}_{0,0}$ are $\mathsf{NTP}_{2}$.
@@
If $M$ is an $^\widehat{\mathcal{L}}$-valued difference field, or a valued difference field in the language $\mathcal{L}_{\sigma}$, as in the proof of Corollary˜3.10 we may expand a suitable pec extension of it to an $\widehat{\mathcal{L}}_{s,\iota,\sigma}$-valued difference field. It follows that it suffices to show that $\mathsf{ACVFA}_{0,0,s,\iota}$ is $\mathsf{NTP}_{2}$.
Suppose for a contradiction that a formula $\varphi(x,y)$ has $\mathsf{TP}_{2}$. We may assume that $x$ is a single variable by Corollary˜4.19. As the theory of $\Gamma$ is $\mathsf{NIP}$ (hence $\mathsf{NTP}_{2}$, see [DGK, Theorem 1.1]) by [DM, Corollary 6.39], by Proposition˜4.37(2) and ˜4.13 the structure $\Gamma_{\mathrm{ind}}(\mathfrak{U})$ is $\mathsf{NTP}_{2}$. Similarly for $k_{\mathrm{ind}}(\mathfrak{U})$, using the fact that $k(\mathfrak{U})\vDash\mathsf{ACFA}$, hence is simple [CHr]. Therefore, by Lemmas˜4.28 and 4.16 we may further assume that $x$ is of sort $\mathrm{VF}$. By Proposition˜4.35 and ˜4.13 we may furthermore assume that $\varphi(x,y)$ is quantifier-free.
Let $(a_{ij})_{i,j<\omega}$ be a strongly indiscernible array witnessing that $\varphi(x,y)$ has $\mathsf{TP}_{2}$. Let $a\vDash\{\varphi(x,a_{i0})\mid i<\omega\}$. By ˜4.4 we may assume that the sequence $(\bar{a}_{i})_{i<\omega}$ is $a$-indiscernible. We run the analogues for the language $\widehat{\mathcal{L}}_{s,\iota,\sigma}$ of steps (I) to (V) in the proof of [CH, Theorem 4.1], but we replace [CH, Lemmas 3.6 and 3.8] by Lemmas˜4.31 and 4.26. We obtain, for $\beta<\omega$, strongly indiscernible arrays $(a_{ij}^{\beta})_{i,j<\omega}$ such that $K\coloneqq\bigcup_{\beta<\omega}a^{\beta}_{00}$ is a substructure of $\mathfrak{U}$ (in fact, a valued difference subfield because, since $\iota$ and $s$ are in the language, the maps $\operatorname{res}$ and $v$ are automatically surjective), the union is increasing, $L\coloneqq K\langle a\rangle$ is an immediate extension of $K$,